# An End-to-End Model for Photo-Sharing Multi-modal Dialogue Generation

### v1

@@
Abstract.
Photo-Sharing Multi-modal dialogue generation requires a dialogue agent not only to generate text responses but also to share photos at the proper moment. Using image text caption as the bridge, a pipeline model integrates an image caption model, a text generation model, and an image generation model to handle this complex multi-modal task. However, representing the images with text captions may loss important visual details and information and cause error propagation in the complex dialogue system. Besides, the pipeline model isolates the three models separately because discrete image text captions hinder end-to-end gradient propagation. We propose the first end-to-end model for photo-sharing multi-modal dialogue generation, which integrates an image perceptron and an image generator with a large language model. The large language model employs the Q-Former to perceive visual images in the input end. For image generation in the output end, we propose a dynamic vocabulary transformation matrix and use straight-through and gumbel-softmax techniques to align the large language model and stable diffusion model and achieve end-to-end gradient propagation. We perform experiments on PhotoChat and DialogCC datasets to evaluate our end-to-end model. Compared with pipeline models, the end-to-end model gains state-of-the-art performances on various metrics of text and image generation. More analysis experiments also verify the effectiveness of the end-to-end model for photo-sharing multi-modal dialogue generation.
@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOMMccs: Computing methodologies Natural language generationccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Image representationsccs: Computing methodologies Reconstruction
1. Introduction
@@
Human-machine dialogue system is a crucial research and application area for the artificial intelligence community (Zhang et al., 2020; Roller et al., 2021; Adiwardana et al., 2020; Yi et al., 2024). With the development of vision-language representation learning, multi-modal dialogue has attracted widespread attention (Das et al., 2017; Mostafazadeh et al., 2017; Shuster et al., 2020; Zheng et al., 2022; Shuster et al., 2021). The multi-modal settings in these literatures are limited to adding visual modality to the input, while the output of the dialogue model remains pure textual modality. Similar to the conversation manner of humans on instant messaging softwares (e.g., Facebook, WhatsApp, and WeChat), an engaging multi-modal dialogue agent should not only perceive and comprehend visual content but also generate images. Therefore, researchers have started to focus on photo-sharing multi-modal dialogue in recent years (Lee et al., 2021; Zang et al., 2021; Sun et al., 2022; Feng et al., 2023; Lee et al., 2024; Aboutalebi et al., 2024), which requires the active behavior of sharing photos in multi-turn dialogue for multi-modal chatbots.
In order to build the photo-sharing multi-modal dialogue system, Sun et al. (2022) propose a pipeline model, Divter, which composes three modules: image caption model, text dialogue model (DialogGPT (Zhang et al., 2020)), and image generation model (DALL-E (Ramesh et al., 2021)). The pipeline model utilizes the image text description as the bridge to connect visual and textual modality. Specifically, for the input end, the text dialogue model perceives photos based on the image descriptions from the image caption model. For the output end, the text dialogue model generates text responses and image descriptions, which are converted into photos by the image generation model.
@@
Although the pipeline model can achieve photo-sharing multi-modal dialogue, the system still faces many challenges. First, the text dialogue model may lose visual details in the photos, because the image caption model will significantly compress visual information during the process of generating image descriptions. As a result, it is difficult for the model to produce elaborate utterance replies. Second, the pipeline model suffers from the severe problem of error propagation. Any small mistake in the image descriptions will be propagated and enlarged into the generated text dialogue responses and images. Third, although the pipeline model includes three neural network models, they are independent and cannot achieve end-to-end gradient propagation and training. So the separate training of these models makes it difficult to capture the internal relationships between different modules, thereby hindering a unified end-to-end photo-sharing multi-modal dialogue model. Besides, the cascaded system has redundant functions and parameters, leading to latency and inefficiency in practical deployment.
Recent progress in multi-modal representation alignment (Radford et al., 2021; Zhang et al., 2024; Bordes et al., 2024), large language models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023) and diffusion models (Ho et al., 2020; Rombach et al., 2022) offer potential solutions and directions for addressing these issues. For one thing, large language models and diffusion models have achieved remarkable results in both text and image generation tasks. Advanced foundation models can improve the final performance of the photo-sharing multi-modal dialogue system. For another thing, current multi-modal large language models integrate visual encoders with large language models, unifying the representation space of vision and text. This enables text generation models to directly perceive the entire image, rather than relying on image captions that may lose visual details. Furthermore, a single multi-modal model performing both image perception and text generation fully liberates the image caption model.
@@
In this work, we propose an end-to-end photo-sharing multi-modal dialogue model to alleviate the challenges in the pipeline model, resorting to recent multi-modal generation models and end-to-end gradient connection techniques (Straight-Through Estimator (Bengio et al., 2013) and Gumbel-Softmax (Jang et al., 2017; Maddison et al., 2017)). First, we use the large language model (Llama 2 (Touvron et al., 2023)) as the backbone of text dialogue generation. Second, Q-Former in BLIP-2 (Li et al., 2023) connects the vision representation encoder (ViT (Dosovitskiy et al., 2021)) with the large language model, which enables the realization of the end-to-end multi-modal model at the input end. Third, the stable diffusion (Rombach et al., 2022) is engaged in image generation in our photo-sharing dialogue system. Especially, we design a dynamic vocabulary transformation matrix to convert LLM vocab into diffusion model vocab, which integrates Straight-Through Estimator (Bengio et al., 2013) and Gumbel-Softmax (Jang et al., 2017; Maddison et al., 2017) to connect the diffusion model with the large language model and realize the end-to-end gradient connection at the output end. Overall, the vision encoder and the diffusion model equip the large language model with the ability to see and draw images at the same time. Q-Former, dynamic vocabulary transformation matrix and Straight-Through Gumbel-Softmax implement end-to-end gradient propagation and joint training.
We conduct experiments on PhotoChat (Zang et al., 2021) and DialogCC (Lee et al., 2024) datasets to verify the effectiveness of our end-to-end photo-sharing multi-modal dialogue model. We employ many diverse textual and visual metrics to evaluate the performances of text and image generation, including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), Inception Score (IS) (Salimans et al., 2016) and Frechet Inception Distance (FID) (Heusel et al., 2017). Our end-to-end model acquires improvements on all metrics compared with the conventional pipeline model (Divter) and LLM-based pipeline model.
Figure 1. Photo-sharing Multi-modal dialogue.
In conclusion, the main contributions can be summarized as follows:
  • We propose the first end-to-end photo-sharing multi-modal dialogue model, which jointly trains the image encoder, large language model and diffusion model to acquire the abilities of image perceiving, text generation and image generation.
  • We design a novel dynamic vocabulary transformation matrix to convert image text description from LLM to diffusion model while retaining gradients.
  • Experimental results on diverse metrics verify the effectiveness of our end-to-end model compared with pipeline models.
2. Model
@@
Figure 2. The overview of our end-to-end photo-sharing multi-modal dialogue generation model, which integrates an image perceptron and an image generator with a large language model. The large language model employs the Q-Former to perceive visual images in the input end. For image generation in the output end, we propose a dynamic vocabulary transformation matrix (DVTM) and use straight-through and gumbel-softmax (STGS) techniques to align the large language model and stable diffusion model and achieve end-to-end gradient propagation.
@@
In this section, we start with the task formalization of photo-sharing multi-modal dialogue (§2.1), then introduce our end-to-end model, which consists of three components: text dialogue generator (§2.2), image perception (§2.3) and image generator (§2.4), end with the description of end-to-end training of our multi-modal dialogue model (§2.5).
2.1. Task Formalization
@@
Given a photo-sharing multi-modal dialogue dataset $\mathcal{D}=\{(U_{i},R_{i})\}_{i}^{I}$, where $\forall~{}i\in\{1,2,\dots,I\}$, $U_{i}$ is the dialogue context, $R_{i}$ is the response regarding to $U_{i}$. Dialogue context $U_{i}$ and response $R_{i}$ could include two modalities: textual elements (e.g., utterances) and visual elements (e.g., images). Concretely, for each dialogue context and response, $U_{i}=\{u_{k}\}_{k}^{K}$ and $R_{i}=\{r_{l}\}_{l}^{L}$ are sequences of multi-modal elements including textual utterances and visual images. $K$ and $L$ are the number of elements in $U_{i}$ and $R_{i}$ respectively. Our goal is to learn an end-to-end photo-sharing multi-modal dialogue model $P(R|U)$ supervised by the dataset $\mathcal{D}$. When given the multi-modal dialogue context $U$, the end-to-end model can generate multi-modal response $R$ including texts and images.
2.2. Text Dialogue Generator
In recent years, large language models have demonstrated powerful text generation ability. Besides, they can also acquire visual ability through extra vision modules, becoming multi-modal large language models. Therefore, we adapt the large language model, Llama (Touvron et al., 2023), as the backbone of our end-to-end photo-sharing multi-modal dialogue model. Specifically, Llama is a sequence-to-sequence generation model based on decoder-only transformer architecture, which autoregressively generates the next token according to the input sequence.
@@ 2.2. Text Dialogue Generator
We input the dialogue context $U$ and response $R$ to the text dialogue generator as the source and target sequence for teacher-forcing training. First, the large language model generates the hidden state representation of the current time step $\bm{h}_{i}$, based on the dialogue context $U$ and the target sequence of the first $i-1$ time steps $r_{\textless i}$. Then, a linear projection layer (LPL) converts the dimension of the hidden state vector $\bm{h}_{i}$ into the size of the large language model vocabulary. This operation can be seen as computing the similarity vector $\bm{o}_{i}$ between $\bm{h}_{i}$ and all tokens in the vocabulary. Finally, we apply softmax to gain the token probability distribution $\bm{p}_{i}$ on the vocabulary of the large language model for the current time step:
(1) $\bm{h}_{i}$ $=\text{LLM}(U,r_{\textless i})$
$\bm{o}_{i}$ $=\text{LPL}(\bm{h}_{i})$
$\bm{p}_{i}$ $=\text{SoftMax}(\bm{o}_{i})$
In the end, we minimize the negative log-likelihood of the target sequence $y$ to optimize the text dialogue generator:
(2) $\mathcal{L}^{\text{t}}=-\text{log}~{}P(R|U)=-\text{log}\prod_{i=1}^{|R|}~{}P(y_{i}|U,r_{\textless i})=-\sum_{i=1}^{|R|}\text{log}~{}\bm{p}_{i}^{r_{i}}$
where $\mathcal{L}^{\text{t}}$ is the notation of text generation loss.
For the inference stage, we sample the token from vocabulary probability distribution $\bm{p}$ and generate the response sequence auto-regressively. Concretely, the decoder-only-based model predicts the next token by encoding the dialogue context and the historical token sequence of the generated response. The text dialogue generator repeats the generation process until the generated response reaches the end token or the max sequence length. Particularly, contents between image start token [IMG] and image end token [/IMG] will be inputted to the image generator, guiding the diffusion model to generate appropriate images. Besides, we also feed the image text description and the corresponding vocabulary probability distribution $\bm{p}_{i}$ to the image generator during the training stage. Our novel gradient connection technique maintains gradient propagation between the large language model and the diffusion model, and hence builds an end-to-end model, which will represented in the following subsection 2.4.
2.3. Image Perceptron
For the photo in the input end, the pipeline model bridges text and vision modality by discrete image text description, which interrupts gradient computation and overlooks visual details. To enable the large language model to directly perceive the visual modality information of the entire image, rather than relying on image descriptions of textual modality, image perceptron attempts to connect a visual encoder to the large language model. There are two main approaches connecting visual encoders and large language models by inserting extra vision-language adaption layers: Q-Former (Li et al., 2023) and linear projection (Liu et al., 2023). We select Q-Former for our end-to-end photo-sharing multi-modal dialogue model, because it encodes images with arbitrary size to a fixed number of visual token representations. This feature is friendly for the multi-modal dialogue task, where users may send photos of various sizes.
Given input image $\bm{I}$, Q-Former generates a fixed number of visual embeddings $\bm{H}^{\text{v}}$ (32 tokens) based on a set number of learnable query embeddings $\bm{Q}$ (32 tokens) and visual features $\widetilde{\bm{H}^{\text{v}}}$ encoded by pretrained ViT encoder (Dosovitskiy et al., 2021):
(3) $\widetilde{\bm{H}^{\text{v}}}$ $=\text{ViT}(\bm{I})$
@@ 2.3. Image Perceptron
After that, visual embeddings $\bm{H}^{\text{v}}$ are fed to the large language model as dialogue context. Since the modality representation conversion occurs entirely in the dense vector space, the image perceptron is an end-to-end model.
2.4. Image Generator
@@
Photo-sharing multi-modal dialogue requires that the multi-modal chatbot can actively share a photo at the right moment, so we design the image generator to be responsible for the image generation function. Diffusion-based models have recently achieved impressive results in image generation tasks (Ramesh et al., 2022; Ho et al., 2020; Rombach et al., 2022). We adapt stable diffusion model (Rombach et al., 2022) in our end-to-end model, as it conducts the diffusion process in latent space not pixel space, offering higher computational efficiency and lower resource requirements. Concretely, stable diffusion consists of a clip text encoder (Radford et al., 2021) and a U-Net model (Ronneberger et al., 2015). The U-Net model iteratively performs the denoising process on the sampled gaussian noise to generate images, guided by the text representations of image text descriptions from the clip text encoder. Similar to the handling of the visual modality in the input end, the pipeline model still drives the vision model through the discrete image text description for image generation in the output end. Our end-to-end model also employs image text description to bridge the large language model and the stable diffusion model, but we propose the Straight-Through Gumbel-Softmax (Bengio et al., 2013; Jang et al., 2017; Maddison et al., 2017) and dynamic vocabulary transformation matrix to implement end-to-end gradient propagation from the image generator to the text dialogue generator.
In this section, we introduce an end-to-end image generator, which ensures gradient flow from the large language model to the stable diffusion model. First, Straight-Through Gumbel-Softmax is presented in section § 2.4.1. Then, we propose a novel dynamic vocabulary transformation matrix in section § 2.4.2, which addresses the sub-word encoding issue of vocabulary in the naive vocabulary transformation matrix. Finally, section § 2.4.3 introduces the stable diffusion model to generate images.
2.4.1. Straight-Through Gumbel-Softmax
The pipeline model completely isolates the text dialogue generation model and the image generation model and only bridges them with the discrete image text caption. For one thing, the textual signal is discrete, interrupting gradient propagation and preventing interaction between text and image generation models. For another thing, the separate training causes the image generation model to be sensitive to the text dialogue generation model. The error accumulation restricts the final performance and user experience of the multi-modal dialogue agent. Based on the straight-through gumbel-softmax (Bengio et al., 2013; Jang et al., 2017; Maddison et al., 2017), we propose an end-to-end photo-sharing multi-modal dialogue model to connect the large language model and stable diffusion model in continuous vector space not discrete textual space. Especially, our end-to-end model contributes to the image generation model not only being robust to the text dialogue model but also using powerful dialogue representations from the large language model.
@@ 2.4.1. Straight-Through Gumbel-Softmax
Concretely, the textual dialogue model applies argmax operation on the vocabulary probability $\bm{p}$ for the next token prediction to obtain the image text caption auto-regressively, which is then delivered to the stable diffusion model to guide photo generation. The argmax and textual input are the sources of weak robustness and the isolation between text dialogue generation and image generation. Compared with the argmax operation that generates deterministic results, gumbel-softmax introduces randomness into the prediction, resulting in a sampled vector. More importantly, gumbel-softmax adapts a reparameterization trick to make the sampling process differentiable.
In particular, we feed the dialogue context $U$ and the partial response sequence of the first $i-1$ time steps $r_{<i}$ into the text dialogue generator, obtaining the token probability distribution vector $\bm{p}_{i}$ of the large language model as shown in Equation 1. Then, we sample noise vector $\bm{g}$ from standard gumbel distribution. Following (Jang et al., 2017; Maddison et al., 2017), the standard gumbel distribution is generated using inverse transform sampling on the random variable $\bm{u}$ from standard uniform distribution:
(4) $\bm{g}$ $\sim Gumbel(0,1)$
$\bm{g}$ $=-\text{log}(-\text{log}(\bm{u}))$
$\bm{u}$ $\sim Uniform(0,1)$
where $Gumbel(0,1)$ and $Uniform(0,1)$ denote standard gumbel distribution and standard uniform distribution, respectively. After that, we add the random noise into the vocabulary probability distribution of the $i$-th time step $\bm{p}_{i}$, and gain a new vocabulary probability distribution $\bm{p}_{i}^{\text{GS}}$ by softmax:
(5) $\bm{p}_{i,j}^{\text{GS}}=\frac{\text{exp}((\text{log}(\bm{p}_{i,j})+\bm{g}_{i,j})/\tau)}{\sum_{k=0}^{|V^{\text{LLM}}|}\text{exp}(\text{log}(\bm{p}_{i,k})+\bm{g}_{i,k})/\tau}$
where $\bm{p}_{i,j}$, $\bm{p}_{i,j}^{\text{GS}}$ and $\bm{g}_{i,j}$ are the $j$-th values of the vectors $\bm{p}_{i}$, $\bm{p}_{i}^{\text{GS}}$ and $\bm{g}_{i}$, respectively. $|V^{\text{LLM}}|$ indicates the size of the large language model vocabulary $V^{\text{LLM}}$. As the softmax temperature $\tau$ approaches 0, samples from the gumbel-softmax distribution become one-hot. Besides, gumbel-softmax conducts a logarithm operation on token probability $\bm{p}_{i}$, which is the result of an exponent arithmetic on similarity vector $\bm{o}_{i}$. Hence, we substitute $\text{log}(\bm{p}_{i})$ with $\bm{o}_{i}$ for computation convenience as follows:
(6) $\bm{p}_{i,j}^{\text{GS}}=\frac{\text{exp}((\bm{o}_{i,j}+\bm{g}_{i,j})/\tau)}{\sum_{k=0}^{|V^{\text{LLM}}|}\text{exp}(\bm{o}_{i,k}+\bm{g}_{i,k})/\tau}$
which is almost equivalent to Equation 5 in math and does not influence model performance in our experiments.
In order to bridge the text dialogue generator and image generator, we still need a discretized one-hot vector to indicate the predicted token in the vocabulary. Consequently, we apply hard gumbel-softmax to avoid the soft categorical probability, which achieves the exact one-hot output $\bm{r}_{i}$ by argmax and retains gradient by straight-through gradient estimator (Bengio et al., 2013) trick:
(7) $\widetilde{\bm{r}_{i}}$ $=\text{argmax}(\bm{p}_{i}^{\text{GS}})$
$\bm{r}_{i}$ $=\widetilde{\bm{r}_{i}}-\text{sg}[\bm{p}_{i}^{\text{GS}}]+\bm{p}_{i}^{\text{GS}}$
@@
where $\text{sg}[\cdot]$ denotes the stop-gradient operation. Straight-through gradient estimator not only ensures exactly discrete vector outputs but also copies gradient from $\bm{r}_{i}$ to $\bm{p}_{i}^{\text{GS}}$. Particularly, since the added noise follows a Gumbel distribution, the distribution of the predicted token from argmax is essentially consistent with the original distribution. Finally, we aggregate one-hot token predictions from the large language model into an image text caption sequence $\bm{R}^{\text{LLM}}=(\bm{r}_{1},\cdots,\bm{r}_{m})$.
2.4.2. Dynamic Vocabulary Transformation Matrix

### v2

@@
Abstract
Photo-Sharing Multi-modal dialogue generation requires a dialogue agent not only to generate text responses but also to share photos at the proper moment. Using image text caption as the bridge, a pipeline model integrates an image caption model, a text generation model, and an image generation model to handle this complex multi-modal task. However, representing the images with text captions may lose important visual details and information and cause error propagation in the complex dialogue system. Besides, the pipeline model isolates the three models separately because discrete image text captions hinder end-to-end gradient propagation. We propose the first end-to-end model for photo-sharing multi-modal dialogue generation, which integrates an image perceptron and an image generator with a large language model. The large language model employs the vision encoder to perceive visual images in the input end. For image generation in the output end, we propose a dynamic vocabulary transformation matrix and use straight-through and gumbel-softmax techniques to align the large language model and stable diffusion model and achieve end-to-end gradient propagation. We perform experiments on PhotoChat and DialogCC datasets to evaluate our end-to-end model. Compared with pipeline models, the end-to-end model gains state-of-the-art performances on various metrics of text and image generation.
Index Terms:
@@ Abstract
I Introduction
@@
Human-machine dialogue system is a crucial research and application area for the artificial intelligence community. With the development of vision-language representation learning, multi-modal human-machine dialogue has attracted widespread attention [1, 2, 3]. However, the multi-modal settings in these literatures and recent multi-modal large language models (MLLMs) (e.g., Llama-Vision [4], Qwen-VL [5]) are limited to adding visual modality to the input, while the output of the dialogue model remains pure textual modality. Similar to the conversation manner of humans on instant messaging softwares (e.g., Facebook, WhatsApp, and WeChat), an engaging multi-modal dialogue agent should not only perceive and comprehend visual content but also generate images. Therefore, researchers have started to focus on photo-sharing multi-modal dialogue in recent years [6, 7, 8, 9, 10, 11], which requires the active behavior of sharing photos in multi-turn dialogue like Figure 1 for multi-modal chatbots.
Figure 1: Photo-sharing multi-modal dialogue.
@@
In order to build the photo-sharing multi-modal dialogue system, [8] proposes a pipeline model, Divter, which composes three modules: image caption model, text dialogue model, and image generation model. The pipeline model utilizes the image text description as the bridge to connect visual and textual modality. Although the pipeline model can achieve photo-sharing multi-modal dialogue, the system still faces many challenges. First, the text dialogue model may lose visual details in the photos, because the image caption model will significantly compress visual information during the process of generating image descriptions. As a result, it is difficult for the model to produce elaborate utterance replies. Second, the pipeline model suffers from the severe problem of error propagation. Any small mistake in the image descriptions will be propagated and enlarged into the generated text dialogue responses and images. Third, the pipeline model includes three neural network models, which are independent and cannot achieve end-to-end gradient propagation and training. So the separate training of these models makes it difficult to capture the internal relationships between different modules. Besides, the cascaded system has redundant functions and parameters, leading to latency and inefficiency in practical deployment.
@@
In this work, we propose an end-to-end photo-sharing multi-modal dialogue model to alleviate the challenges in the pipeline model, resorting to recent MLLMs, diffusion models [12] and end-to-end gradient connection techniques (Straight-Through Estimator [13] and Gumbel-Softmax [14, 15]). For one thing, we use the MLLM (i.e., Llama-Vision) as the backbone of dialogue generation, which enables the realization of the end-to-end multi-modal model at the input end. For another thing, the stable diffusion [12] is engaged in image generation. Especially, inspired by [16] and [17], we design a dynamic vocabulary transformation matrix to convert LLM vocabulary into diffusion model vocabulary, which integrates Straight-Through Estimator [13] and Gumbel-Softmax [14, 15] to connect the diffusion model with the MLLM and realize the end-to-end gradient connection at the output end. Overall, the MLLM and the diffusion model equip the large language model with the ability to see and draw images at the same time.
In conclusion, the main contributions can be summarized as follows: i) We propose the first end-to-end photo-sharing multi-modal dialogue model, which jointly trains the MLLM and diffusion model to acquire the abilities of image perceiving, text generation and image generation. ii) We design a novel dynamic vocabulary transformation matrix and apply straight-through and gumbel-softmax techniques to convert image text description from LLM to diffusion model while retaining gradients. iii) Experimental results on diverse metrics verify the effectiveness of our end-to-end model compared with pipeline models  [^1] We public our code at https://github.com/guopeiming/E2E_PSDG..
II Related Work
With the development of vision-language representation learning in this decade, multi-modal dialogue has attracted widespread attention [1, 2, 3]. However, conventional multi-modal dialogue tasks consider vision modality only at the input end. Therefore, researchers have focused on photo-sharing multi-modal dialogue in recent years [6, 7, 8, 9, 10, 11]. [7] proposes a retrieval model to predict the image shared in the dialogue, which computes the similarity between dialogue context and image. [9] also models photo-sharing multi-modal dialogue by a retrieval task. Different from only predicting images in [7], [9] retrieves suitable textual utterances and shared photos for conversations. In addition to the retrieval model, [8] presents a generative model, Divter, for photo-sharing multi-modal dialogue generation. Specifically, Divter adapts a text dialogue model as the backbone, and perceives photos based on discrete image descriptions from an image caption model for the input end. For the output end, the text dialogue model generates text responses and image descriptions, which are converted into photos by an image generation model.
Another important related work for this paper is multi-modal generation models, including image generation models and MLLMs. Image generation models have gained great attention in the deep learning community in this decade. In the beginning, researchers employ VAE and GAN models to generate image examples. Following, VQ-VAE [18] and VQ-GAN [19] quantize latent vectors into discrete codebook to improve VAE and GAN models. Recently, diffusion-based models have shown better image generation ability [20], like DALL-E-2 [21] and stable diffusion [12]. Most MLLMs [22, 5] integrate a vision encoder into LLMs, which learn a mapping between text modality (large language model) and vision modality (vision encoder). However, these MLLMs acquire multi-modal capability only at the input end, employing linear projection [22], Q-Former [5] or cross-attention as [4] as the modality mapping adapter. A few MLLMs [23, 24] have recently implemented visual ability at both the input and output end by vision tokenizer.
III Model
@@
Figure 2: The overview of our end-to-end photo-sharing multi-modal dialogue generation model, which integrates an image perceptron and an image generator with a large language model. The large language model employs the image perceptron to perceive visual images in the input end. For image generation in the output end, we propose a dynamic vocabulary transformation matrix (DVTM) and use straight-through and gumbel-softmax (STGS) techniques to align the large language model and stable diffusion model and achieve end-to-end gradient propagation.
@@
In this section, we start with the task formalization of photo-sharing multi-modal dialogue (§III-A), then introduce our end-to-end model shown in Figure 2, which consists of three components: text dialogue generator (§III-B), image perception (§III-C) and image generator (§III-D), end with the description of end-to-end training (§III-E).
III-A Task Formalization
@@
Given a photo-sharing multi-modal dialogue dataset $\mathcal{D}=\{(U_{i},R_{i})\}_{i}^{I}$, where $\forall~{}i\in\{1,2,\dots,I\}$, $U_{i}$ is the dialogue context, $R_{i}$ is the response regarding to $U_{i}$. Concretely, for each dialogue context and response, $U_{i}=\{u_{k}\}_{k}^{K}$ and $R_{i}=\{r_{l}\}_{l}^{L}$ are sequences of multi-modal elements including textual utterances and visual images. $K$ and $L$ are the number of elements in $U_{i}$ and $R_{i}$ respectively. Our goal is to learn an end-to-end photo-sharing multi-modal dialogue model $P(R|U)$ supervised by the dataset $\mathcal{D}$. When given the multi-modal dialogue context $U$, the end-to-end model can generate multi-modal response $R$ including texts and images.
III-B Text Dialogue Generator
@@
We input the dialogue context $U$ and response $R$ to the text dialogue generator as the source and target sequence for teacher-forcing training on the vocabulary probability distribution $\bm{p}_{i}$ of the $i$-th time step:
$\bm{p}_{i}$ $=\text{LLM}(U,r_{\textless i})$ (1)
$\mathcal{L}^{\text{t}}$ $=-\text{log}~{}P(R|U)=-\Sigma_{i=1}^{|R|}\text{log}~{}\bm{p}_{i}^{r_{i}},$
where $\bm{p}_{i}^{r_{i}}$ is probability of the token $r_{i}$. For the inference stage, LLM auto-regressively generates textual utterances or image captions. Particularly, the image caption and the corresponding vocabulary probability distribution $\bm{p}_{i}$ will be inputted to the image generator, which employs the diffusion model and our novel gradient connection technique to generate images and maintain gradient propagation.
III-C Image Perceptron
@@
III-D Image Generator
@@
Photo-sharing multi-modal dialogue requires the multi-modal chatbot to share a photo at the right moment actively, so we design the image generator to be responsible for the image generation function. We adapt the stable diffusion model [12] in our end-to-end system for its impressive generation performance and computational efficiency. Similar to the handling of the visual modality in the input end, the pipeline model still drives the vision model through the discrete image text description for image generation in the output end. Our end-to-end model also employs image text description to bridge the large language model and the stable diffusion model, but we propose the Straight-Through Gumbel-Softmax [13, 14, 15] and dynamic vocabulary transformation matrix to implement end-to-end gradient propagation from the image generator to the text dialogue generator.
III-D1 Straight-Through Gumbel-Softmax
@@
The pipeline model completely isolates the text dialogue generation model and the image generation model and only bridges them with the discrete image text caption. Based on the straight-through gumbel-softmax [13, 14, 15], we propose an end-to-end model to connect the LLM and stable diffusion model in continuous vector space not discrete textual space. Concretely, the text dialogue generator applies argmax operation to generate an image caption, which is then delivered to the stable diffusion model to guide photo generation. The argmax and textual input are the sources of weak robustness and the isolation between text dialogue generation and image generation. Compared with the argmax operation, gumbel-softmax introduces randomness into the prediction, resulting in a sampled vector. More importantly, gumbel-softmax adapts a reparameterization trick to make the sampling process differentiable. Specifically, we first sample random variable $\bm{g}$ and add it into the vocabulary probability distribution $\bm{p}_{i}$ to gain a new one $\bm{p}_{i}^{\text{GS}}$. Then we apply hard gumbel-softmax to avoid the soft categorical probability, which achieves the exact one-hot output $\bm{r}_{i}$ by argmax and retains gradient by straight-through gradient estimator [13]:
$\bm{g}$ $\sim Gumbel(0,1)$ (2)
$\bm{p}_{i,j}^{\text{GS}}$ $=\frac{\text{exp}((\text{log}(\bm{p}_{i,j})+\bm{g}_{i,j})/\tau)}{\sum_{k=0}^{|V^{\text{LLM}}|}\text{exp}(\text{log}(\bm{p}_{i,k})+\bm{g}_{i,k})/\tau}$
$\bm{r}_{i}$ $=\text{argmax}(\bm{p}_{i}^{\text{GS}})-\text{sg}[\bm{p}_{i}^{\text{GS}}]+\bm{p}_{i}^{\text{GS}}.$
@@
where $\bm{p}_{i,j}$, $\bm{p}_{i,j}^{\text{GS}}$ and $\bm{g}_{i,j}$ are the $j$-th values of the vectors $\bm{p}_{i}$, $\bm{p}_{i}^{\text{GS}}$ and $\bm{g}_{i}$, respectively. $|V^{\text{LLM}}|$ indicates the size of the LLM vocabulary $V^{\text{LLM}}$. As the temperature $\tau$ approaches 0, samples from the gumbel-softmax distribution become one-hot. $\text{sg}[\cdot]$ is the stop-gradient operation for straight-through gradient estimator, which not only ensures exactly discrete vector outputs but also copies gradient from $\bm{r}_{i}$ to $\bm{p}_{i}^{\text{GS}}$. Finally, we aggregate one-hot token predictions from the large language model into an image text caption sequence $\bm{R}^{\text{LLM}}=(\bm{r}_{1},\cdots,\bm{r}_{m})$.
III-D2 Dynamic Vocabulary Transformation Matrix